# A stochastic Gordon-Loeb model for optimal cybersecurity investment under clustered attacks

### v1

@@ 4.1 Specification of the Model Parameters
We employ a security breach probability function $S$ of class I, as defined in equation (2). The parameters values in Table I are consistent with those in [27], and also in line with previous works (see [3] and [28]). Taking $v,a,b$ as in Table I, a plot of the function $h\mapsto S_{I}(h,v)$ is shown in Figure 2.
@@
The parameters of the stochastic intensity of the Hawkes process (see Table II) are chosen to generate on average approximately 60 cyberattacks per year. We believe that this is a reasonable figure, in the absence of reliable estimates of the number of cyberattacks targeting a single entity. [^7] Empirical estimates of the intensity of cyberattacks can be found in the recent works [19], [20], [29]). However, these estimates are not suitable for our purposes, since they are based on the number of attacks at a worldwide scale, while our model takes the viewpoint of a single entity.
@@
[…]
@@ 4.2 Numerical Solution of the HJB-PIDE
As shown in Section 3, determining the optimal cybersecurity investment requires the solution of the non-linear PIDE (16). Due to the complexity of the problem, one cannot expect to find explicitly an analytical solution and, hence, numerical methods are required. We opt for the method of lines, as described in [30]. This technique consists in discretizing the PIDE in the spatial domain $(\lambda,h)\in(0,\infty)\times\mathbb{R}_{+}$ but not in time, and then in integrating the semi-discrete problem as a system of ODEs. In our setting, we discretize the $(\lambda,h)$ dimensions with a central difference and then numerically solve the resulting ODE system. Similarly to the case of PIDEs arising in Lévy models (see, e.g., [31]), the unbounded space domain $(0,\infty)\times\mathbb{R}_{+}$ is localized into a bounded domain $[\lambda_{\text{min}},\lambda_{\text{max}}]\times[h_{\text{min}},h_{\text{max}}]$. We refer to Algorithm 1 for a precise description of the implementation of this method.
@@
The value $h_{\text{min}}=0$ corresponds to the absence of cybersecurity investment, while $h_{\text{max}}=50$ represents an upper bound which is rarely achieved in our setup under the standard parameter set. We choose $\lambda_{\text{min}}=\lambda_{0}$, which coincides with the lower bound of the stochastic intensity $\lambda_{t}$, see equation (19). We set $\lambda_{\text{max}}=\mathbb{E}[\lambda_{T}]+7\sqrt{\text{Var}[\lambda_{T}]}\approx 216.6$, in order to ensure that the truncation of the intensity domain does not have any material impact on our numerical results. The value function $V$ is extrapolated beyond $[\lambda_{\text{min}},\lambda_{\text{max}}]$ by setting
@@
analogously to the scheme implemented in [32, Section 5.1]. When plotting the function $V$ in Section 5, we shall consider a subinterval of $[\lambda_{\text{min}},\lambda_{\text{max}}]$: intensity values close to $\lambda_{\text{max}}$ are rarely achieved and might lead to numerical instabilities.
@@
2:Discretize $[\lambda_{\text{min}},\lambda_{\text{max}}]$, with $\lambda_{0}=\lambda_{\text{min}},\lambda_{N}=\lambda_{\text{max}}$ and $\lambda_{n}-\lambda_{n-1}=\Delta\lambda$, for $n=1,\ldots,N$.
3:Discretize $[h_{\text{min}},h_{\text{max}}]$, with $h_{0}=h_{\text{min}},h_{M}=h_{\text{max}}$ and $h_{m}-h_{m-1}=\Delta h$, for $m=1,\ldots,M$.
@@
8:Solve the ODE system given for all $n,m$ by {dmath*} V’_n,m(t)=ξ(λ_n-α)Vn+1,m(t)-Vn-1,m(t)2Δλ+ρh Vn,m+1(t)-Vn,m-1(t)2Δh - λ_n(V_n+~n∧N,m(t)-V_n,m(t)) - ¯η(v-S(h_m,v))λ_n - ((Vn,m+1(t)-Vn,m-1(t)2Δh-1)+)22γ, $V_{n,m}(T)=U(h_{m}).$
We have implemented Algorithm 1 in Python, using the built-in ODE solver scipy.integrate.solve_ivp. We make use of an implicit Runge-Kutta method of the Radau IIA family of order 5 (see [33] for further details).
@@ 4.3 Optimal Cybersecurity Investment Rate
2:Discretize $[t_{\text{min}},t_{\text{max}}]$, with $t_{0}=t_{\text{min}},t_{I}=t_{\text{max}}$ and $t_{i}-t_{i-1}=\Delta t$, for $i=1,\ldots,I$.
3:Discretize $[\lambda_{\text{min}},\lambda_{\text{max}}]$, with $\lambda_{0}=\lambda_{\text{min}},\lambda_{N}=\lambda_{\text{max}}$ and $\lambda_{n}-\lambda_{n-1}=\Delta\lambda$, for $n=1,\ldots,N$.
4:Discretize $[h_{\text{min}},h_{\text{max}}]$, with $h_{0}=h_{\text{min}},h_{M}=h_{\text{max}}$ and $h_{m}-h_{m-1}=\Delta h$, for $m=1,\ldots,M$.
@@ 5.3 Comparison with a Static Investment Strategy
The optimal investment rate $z^{*}$ characterized in Theorem 1 represents the real-time adaptive cybersecurity policy that best responds to the arrival of cyberattacks. In order to assess whether the adoption of an adaptive dynamic strategy provides a tangible benefit, we compare it against the best constant investment strategy, i.e., the strategy $z_{t}=\bar{z}$, for all $t\in[0,T]$, that maximizes the benefit-cost trade-off functional $J$. When investing according to a costant rate $\bar{z}$, the benefit-cost trade-off functional $J$ takes the following form:
$J(t,\lambda,h;\bar{z})$ $=\int_{t}^{T}\bar{\eta}\bigl{(}v-S(H_{s}^{t,h,\bar{z}},v)\bigr{)}\mathbb{E}[\lambda_{s}^{t,\lambda}]\mathrm{d}s$ (20)
$\quad-(T-t)(\bar{z}+\frac{\gamma}{2}\bar{z}^{2})+U(H_{T}^{t,h,\bar{z}}),$
@@
%$\%\text{gain}(t,\lambda,h):=100\times\frac{V(t,\lambda,h)-J(t,\lambda,h;\bar{z}^{*})}{J(t,\lambda,h;\bar{z}^{*})}.$ (22)
@@ 5.4 Comparison with a Standard Poisson Model
$\frac{\partial V^{P}}{\partial t}-\rho h\frac{\partial V^{P}}{\partial h}+\lambda^{P}\bar{\eta}\,(v-S(h,v))+\frac{\bigl{(}(\frac{\partial V^{P}}{\partial h}-1)^{+}\bigr{)}^{2}}{\gamma}=0,$
@@
    $\lambda^{P}_{e}=\frac{\lambda_{0}\xi}{\xi-\beta}+\frac{1-e^{-\xi T}}{T(\xi-\beta)}\left(\lambda_{0}-\frac{\lambda_{0}\xi}{\xi-\beta}\right)\approx 60.77.$ (26)
@@ 5.4.1 Value Functions and Optimal Cybersecurity Policies
(a) Value functions $V(t,\lambda^{P}_{b},h)$ and $V^{P}_{b}(t,h)$
@@
Figure 8: Comparison with a Poisson model with constant intensity $\lambda^{P}_{e}=60.77$.
@@ 5.4.2 Relative Gain
$J(t,\lambda,h;z^{P*})$ $=\int_{t}^{T}\left(\bar{\eta}(v-S(H_{s}^{t,h,z^{P*}},v))\mathbb{E}[\lambda_{s}^{t,\lambda}]\right)\mathrm{d}s$
$\;-\int_{t}^{T}\left(z^{P*}_{s}+\frac{\gamma}{2}(z^{P*}_{s})^{2}\right)\mathrm{d}s+U\bigl{(}H_{T}^{t,h,z^{P*}}\bigr{)},$
where $H^{t,h,z^{P*}}_{s}$ is defined as in (14) with $z=z^{P*}$ and $\mathbb{E}[\lambda_{s}^{t,\lambda}]$ can be computed explicitly by Proposition 1. Once the PDE (23) is numerically solved, $z^{P*}$ can be computed via Algorithm 2 taking $\lambda$ constant. Our numerical implementation of the Poisson-based model adopts the following specification:
@@
%$\%\text{gain}^{P}(t,\lambda,h):=100\times\frac{V(t,\lambda,h)-J(t,\lambda,h;z^{P*})}{J(t,\lambda,h;z^{P*})}.$ (27)
Figure 9 reports the quantity %$\%\text{gain}^{P}(t,\lambda,h)$ comparing the Hawkes-based model against two Poisson-based benchmarks with intensities $\lambda^{P}_{b}$ and $\lambda^{P}_{e}$, as considered above. For $h=0$, the gain increases with $\lambda$, ranging between $7.6\%$ and $11.4\%$ for $\lambda^{P}_{b}$, and between $0.5\%$ and $0.95\%$ for $\lambda^{P}_{e}$. For $h=20$, the gain becomes nearly constant in $\lambda$. The fact that the gain for $\lambda^{P}_{e}$ is limited can be explained by the fact that the objective functional (10) is linear with respect to the losses and $\lambda^{P}_{e}$ is chosen in such a way that the Poisson-based model generates the same expected losses of our Hawkes-based model.

### v2

@@ 4.1 Specification of the Model Parameters
We employ a security breach probability function $S$ of class I, as defined in (2). The parameters values in Table I are consistent with those determined in [47] through a calibration based on synthetic data and also used as benchmark inputs in subsequent works (see, e.g., [55, 43]). Taking $v,a,b$ as in Table I, the function $h\mapsto S_{I}(h,v)$ is plotted in Figure 2.
@@
The parameters of the stochastic intensity of the Hawkes process (see Table II) are chosen to generate on average approximately 60 cyberattacks per year. We believe that this is a reasonable figure, in the absence of reliable estimates of the number of cyberattacks targeting a single entity. [^7] Empirical estimates of the intensity of cyberattacks can be found in the recent works [6], [11], [38]. However, these estimates are not suitable for our purposes, since they are based on the number of attacks at a worldwide scale, while our model takes the viewpoint of a single entity. An average of 60 cyberattacks per year is in line with the number of attacks per organisation reported in the Cyber Readiness Report 2024 by Hiscox (see [32]).
@@
We consider a one-year planning horizon ($T=1$) and set an average loss of $10\text{k}\mathdollar$ for each successful breach, resulting in a total expected annual loss of approximately $390\text{k}\mathdollar$ without cybersecurity investments, which is in the same order of magnitude of [55]. The depreciation rate is set at $\rho=0.2$, consistently with the technological depreciation rates considered in [35]. The parameter $\gamma$ is set at a rather low value, in order to avoid an excessive penalization of large investment rates. Finally, we choose $U(h)=\sqrt{h}$, representing a strictly increasing and concave CRRA utility function.
@@ 4.2 Numerical Solution of the HJB-PIDE
As shown in Section 3, determining the optimal cybersecurity investment requires the solution of the non-linear PIDE (16). Due to the complexity of the problem, one cannot expect to find explicitly an analytical solution and, hence, numerical methods are required. We opt for the method of lines, as described in [57]. This technique consists in discretizing the PIDE in the spatial domain $(\lambda,h)\in(0,\infty)\times\mathbb{R}_{+}$ but not in time, and then in integrating the semi-discrete problem as a system of ODEs. In our setting, we discretize the $(\lambda,h)$ dimensions with a central difference and then numerically solve the resulting ODE system. Similarly to the case of PIDEs arising in Lévy models (see, e.g., [14]), the unbounded space domain $(0,\infty)\times\mathbb{R}_{+}$ is localized into a bounded domain $\lambda_{\text{min}},\lambda_{\text{max}}]\times[h_{\text{min}},h_{\text{max}}]$. We refer to Algorithm 1 for a precise description of the implementation of this method.
@@
The value $h_{\text{min}}=0$ corresponds to the absence of cybersecurity investment, while $h_{\text{max}}=50$ represents an upper bound which is rarely achieved in our setup under the standard parameter set. We choose $\lambda_{\text{min}}=\lambda_{0}$, which coincides with the lower bound of the stochastic intensity $\lambda_{t}$, see equation (19). We set $\lambda_{\text{max}}=\mathbb{E}[\lambda_{T}]+7\sqrt{\text{Var}(\lambda_{T})}\approx 216$, in order to ensure that the truncation of the intensity domain does not have any material impact on our numerical results. The value function $V$ is extrapolated beyond $\lambda_{\text{min}},\lambda_{\text{max}}]$ by setting
@@
analogously to the scheme implemented in [41, Section 5.1]. When plotting the function $V$ in Section 5, we shall consider a subinterval of $\lambda_{\text{min}},\lambda_{\text{max}}]$: intensity values close to $\lambda_{\text{max}}$ are rarely achieved and might lead to numerical instabilities.
@@
2:Discretize $\lambda_{\text{min}},\lambda_{\text{max}}]$, with $\lambda_{0}=\lambda_{\text{min}},\lambda_{N}=\lambda_{\text{max}}$ and $\lambda_{n}-\lambda_{n-1}=\Delta\lambda$, for $n=1,\ldots,N$.
3:Discretize $h_{\text{min}},h_{\text{max}}]$, with $h_{0}=h_{\text{min}},h_{M}=h_{\text{max}}$ and $h_{m}-h_{m-1}=\Delta h$, for $m=1,\ldots,M$.
@@
8:Solve the ODE system given for all $n,m$ by
$V^{\prime}_{n,m}(t)=\xi(\lambda_{n}-\alpha)\frac{V_{n+1,m}(t)-V_{n-1,m}(t)}{2\Delta\lambda}+\rho h{\frac{V_{n,m+1}(t)-V_{n,m-1}(t)}{2\Delta h}}-\lambda_{n}(V_{n+\tilde{n}\wedge N,m}(t)-V_{n,m}(t))-\bar{\eta}(v-S(h_{m},v))\lambda_{n}-\frac{\left(\left(\frac{V_{n,m+1}(t)-V_{n,m-1}(t)}{2\Delta h}-1\right)^{+}\right)^{2}}{2\gamma},$
$V_{n,m}(T)=U(h_{m}).$
We have implemented Algorithm 1 in Python, using the built-in ODE solver scipy.integrate.solve_ivp. We make use of an implicit Runge-Kutta method of the Radau IIA family of order 5 (see [25] for further details). The computations were performed on an Intel Xeon E5520 CPU equipped with 32 GB of RAM, requiring approximately 23 hours to compute the value function and the corresponding optimal control. To empirically assess the convergence of Algorithm 1, we verified that standard discretization error metrics decrease consistently as the spatial grid in $(\lambda,h)$ is refined.
@@ 4.3 Optimal Cybersecurity Investment Rate
2:Discretize $t_{\text{min}},t_{\text{max}}]$, with $t_{0}=t_{\text{min}},t_{I}=t_{\text{max}}$ and $t_{i}-t_{i-1}=\Delta t$, for $i=1,\ldots,I$.
3:Discretize $\lambda_{\text{min}},\lambda_{\text{max}}]$, with $\lambda_{0}=\lambda_{\text{min}},\lambda_{N}=\lambda_{\text{max}}$ and $\lambda_{n}-\lambda_{n-1}=\Delta\lambda$, for $n=1,\ldots,N$.
4:Discretize $h_{\text{min}},h_{\text{max}}]$, with $h_{0}=h_{\text{min}},h_{M}=h_{\text{max}}$ and $h_{m}-h_{m-1}=\Delta h$, for $m=1,\ldots,M$.
@@ 5.3 Comparison with a Static Investment Strategy
The optimal investment rate $z^{*}$ characterized in Theorem 1 represents the real-time adaptive cybersecurity policy that best responds to the arrival of cyberattacks.
In order to assess whether the adoption of an adaptive dynamic strategy provides a tangible benefit, we compare it against the best constant investment strategy, i.e., the strategy $z_{t}=\bar{z}$, for all $t\in[0,T]$, that maximizes the benefit-cost trade-off functional $J$. When investing according to a costant rate $\bar{z}$, the benefit-cost trade-off functional $J$ takes the following form:
$J(t,\lambda,h;\bar{z})$ $=\int_{t}^{T}\bar{\eta}\bigl(v-S(H_{s}^{t,h;\bar{z}},v)\bigr)\mathbb{E}[\lambda_{s}^{t,\lambda}]\mathrm{d}s$ (20)
$\quad-(T-t)(\bar{z}+\frac{\gamma}{2}\bar{z}^{2})+U(H_{T}^{t,h;\bar{z}}),$
@@
$\%\text{gain}(t,\lambda,h):=100\times\frac{V(t,\lambda,h)-J(t,\lambda,h;\bar{z}^{*})}{J(t,\lambda,h;\bar{z}^{*})}.$ (22)
@@ 5.4 Comparison with a Standard Poisson Model
$\frac{\partial V^{P}}{\partial t}-\rho h\frac{\partial V^{P}}{\partial h}+\lambda^{P}\bar{\eta}\,(v-S(h,v))+\frac{\left((\frac{\partial V^{P}}{\partial h}-1)^{+}\right)^{2}}{2\gamma}=0,$
@@
    $\lambda^{P}_{e}=\frac{\lambda_{0}\xi}{\xi-\beta}+\frac{1-e^{-\xi T}}{T(\xi-\beta)}\left(\lambda_{0}-\frac{\lambda_{0}\xi}{\xi-\beta}\right)\approx 61.$ (26)
@@ 5.4.1 Value Functions and Optimal Cybersecurity Policies
(a) Value functions $V(t,\lambda^{P}_{b},h)$ and $V^{P}_{b}(t,h)$ .
@@
Figure 8: Comparison with a Poisson model with constant intensity $\lambda^{P}_{e}=61$.
@@ 5.4.2 Relative Gain
$J(t,\lambda,h;z^{P*})$ $=\int_{t}^{T}\left(\bar{\eta}(v-S(H_{s}^{t,h;z^{P*}},v))\mathbb{E}[\lambda_{s}^{t,\lambda}]\right)\mathrm{d}s$
$\;-\int_{t}^{T}\left(z^{P*}_{s}+\frac{\gamma}{2}(z^{P*}_{s})^{2}\right)\mathrm{d}s+U\bigl(H_{T}^{t,h;z^{P*}}\bigr),$
where $H^{t,h;z^{P*}}_{s}$ is defined as in (14) with $z=z^{P*}$ and $\mathbb{E}[\lambda_{s}^{t,\lambda}]$ can be computed explicitly by Proposition 1. Once the PDE (23) is numerically solved, $z^{P*}$ can be computed via Algorithm 2 taking $\lambda$ constant. Our numerical implementation of the Poisson-based model adopts the following specification:
@@
$\%\text{gain}^{P}(t,\lambda,h):=100\times\frac{V(t,\lambda,h)-J(t,\lambda,h;z^{P*})}{J(t,\lambda,h;z^{P*})}.$ (27)
Figure 9 reports the quantity $\%\text{gain}^{P}(t,\lambda,h)$ comparing the Hawkes-based model against two Poisson-based benchmarks with intensities $\lambda^{P}_{b}$ and $\lambda^{P}_{e}$, as considered above. For $h=0$, the gain increases with $\lambda$, ranging between $7.6\%$ and $11.4\%$ for $\lambda^{P}_{b}$, and between $0.04\%$ and $0.6\%$ for $\lambda^{P}_{e}$. For $h=20$, the gain becomes nearly constant in $\lambda$. The fact that the gain for $\lambda^{P}_{e}$ is limited can be explained by the fact that the objective functional (10) is linear with respect to the losses and $\lambda^{P}_{e}$ is chosen in such a way that the Poisson-based model generates the same expected losses of our Hawkes-based model.
@@ 5.4.3 Adaptive Dynamics of the Optimal Investment Policy
6 Implications for Cyber-insurance
In this section, we connect our dynamic model for optimal cybersecurity investment to actuarial aspects of cyber-insurance. From an insurance viewpoint, the cybersecurity investment policy acts as a form of self-insurance (see [10]): by investing in cybersecurity, the decision maker changes the distribution of aggregate losses over the planning horizon. Cybersecurity investment is therefore a form of prevention, which represents a crucial aspect of cyber-insurance, as discussed in Section 6.1 below. In turn, prevention measures such as cybersecurity investment have an impact on the determination of cyber-insurance premia, as shown in Section 6.2 below.
6.1 The Role of Prevention in Cyber-insurance
Cyber-insurance contracts increasingly go beyond the pure compensation of financial losses and often include prevention requirements and bundled cyber-risk mitigation services. In this sense, cyber-insurance combines risk transfer with risk mitigation services and incentives, promoting best practices in cyber-risk management (see [40] for a recent analysis).
In our model, prevention is represented by a cybersecurity investment policy $z\in\mathcal{Z}$ and has a precise actuarial meaning: it changes the aggregate losses $L_{T}^{z}$ over the planning horizon by reducing the breach probability at attack times from the baseline $v$ to $S(H_{\tau_{i}}^{z},v)$. Note that, by lowering the breach probability at attack times, in our model prevention does not only reduce the expected loss, but also reduces loss variability. As a result, any underwriting or pricing rule that depends on the level of risk should, in principle, depend on the prevention policy (or on observable proxies of it, as discussed below). Our framework allows quantifying the effect of optimal prevention, which is an essential input for the insurer in order to determine cybersecurity requirements, incentive schemes, and premium differentiation rules across policyholders (see Section 6.2).
A key practical issue is that prevention is only partially observable, creating asymmetric information and moral hazard. Two implementation channels discussed in the literature are consistent with our framework. First, prevention may be delegated to the insurer or to an IT security provider in partnership with the insurer, which is particularly relevant for small firms or policyholders with limited internal expertise. This scheme makes prevention partly contractible and allows joint pricing of risk transfer together with mitigation services (see [59]). Second, when prevention is handled internally by the entity (as a self-protection measure), insurers may rely on incentive-based contracting. One approach is to condition the premium on minimum protection measures, combined with audits and penalties if requirements are not met (see [33, 37]). Another approach is to rely on principal–agent incentive schemes, as considered in [12, 42], linking part of the compensation to measurable reduction in risk or losses. In both cases, the insurer needs quantitative tools to assess the impact of prevention measures, as considered in our study.
6.2 Implications for Premium Determination
To quantify the impact of prevention on cyber-insurance premia, we apply the standard deviation principle (see, e.g., [53]), for an arbitrary cybersecurity investment policy $z\in\mathcal{Z}$:
$\pi(L^{z}_{T}):=\mathbb{E}[L^{z}_{T}]+\theta\,\sigma(L^{z}_{T}),$ (28)
where $\mathbb{E}[L_{T}^{z}]$ is the pure premium and $\theta\sigma(L_{T}^{z})$ is a loading proportional to the standard deviation $\sigma(L^{z}_{T})$ of $L^{z}_{T}$, with $\theta$ typically in $[0.1,0.5]$. This principle is stable with respect to the monetary unit and captures the impact of prevention on both the expected loss and the dispersion of losses.
We first compare the expected aggregate loss without additional investment, $\mathbb{E}[L_{T}^{0}]$, with the expected aggregate loss under the optimal prevention policy $z^{\ast}$, denoted $\mathbb{E}[L_{T}^{z^{\ast}}]$, corresponding to (28) with $z=z^{*}$ and $\theta=0$. We compute $\mathbb{E}[L_{T}^{0}]$ by Proposition 2, while $\mathbb{E}[L_{T}^{z^{\ast}}]$ is estimated by Monte Carlo over 100,000100{,}000 samples, where $z^{\ast}$ is computed via Algorithm 2 starting from $h=0$. Under the standard parameter set, we obtain (in k\$)
$\mathbb{E}[L_{T}^{0}]=394.98\qquad\text{and}\qquad\mathbb{E}[L_{T}^{z^{\ast}}]=141.77,$
corresponding to a reduction of approximately $65\%$ in the pure premium under the optimal dynamic prevention strategy.
The variance of the losses $L^{z}_{T}$ associated to a given cybersecurity investment policy $z\in\mathcal{Z}$ is given in the following lemma. We denote $\bar{\eta}:=\mathbb{E}[\eta_{i}]$ and $\sigma^{2}_{\eta}:=\mathrm{Var}(\eta_{i})$, for $i\in\mathbb{N}^{*}$.
Lemma 1.
Suppose that Assumptions (B) and (C) hold. Assume furthermore that $\eta_{i}\in L^{2}(\mathbb{P})$, for all $i\in\mathbb{N}^{*}$. Then, for every $z\in\mathcal{Z}$, it holds that
$\mathrm{Var}(L^{z}_{T})$
$\;=\mathbb{E}\left[\int_{0}^{T}\Bigl(\sigma^{2}_{\eta}\,S(H^{z}_{t},v)+\bar{\eta}^{2}S(H^{z}_{t},v)\bigl(1-S(H^{z}_{t},v)\bigr)\Bigr)\lambda_{t}\mathrm{d}t\right]$
$\quad+\bar{\eta}^{2}\,\mathrm{Var}\left(\int_{0}^{T}S(H^{z}_{t},v)\,\mathrm{d}N_{t}\right).$
Proof.
Let $L_{T}^{z}$ be defined as in (8), where $z=(z_{t})_{t\in[0,T]}$ is an admissible cybersecurity investment process. By the law of total variance (see, e.g., [9, Problem 34.10]), we can compute
$\mathrm{Var}(L_{T}^{z})$ $=\mathbb{E}\left[\mathrm{Var}\left(\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Bigg|\mathcal{F}_{T}\right)\right]$ (29)
$\quad+\mathrm{Var}\left(\mathbb{E}\left[\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Bigg|\mathcal{F}_{T}\right]\right)$
For the first term on the right-hand side in (29) it holds that
$\mathrm{Var}\left(\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Bigg|\mathcal{F}_{T}\right)=\sum_{i=1}^{N_{T}}\mathrm{Var}\left(\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Big|\mathcal{F}_{T}\right)$
$=\sum_{i=1}^{N_{T}}\Bigl(\sigma^{2}_{\eta}\,S(H^{z}_{\tau_{i}},v)+\bar{\eta}^{2}S(H^{z}_{\tau_{i}},v)\bigl(1-S(H^{z}_{\tau_{i}},v)\bigr)\Bigr),$
while for the second term we have that
$\mathbb{E}\left[\sum_{i=1}^{N_{T}}\eta_{i}B_{i}^{S(H^{z}_{\tau_{i}},v)}\Bigg|\mathcal{F}_{T}\right]=\bar{\eta}\sum_{i=1}^{N_{T}}S(H^{z}_{\tau_{i}},v).$
Therefore, we can write that
$\mathrm{Var}(L^{z}_{T})$ $=\sigma^{2}_{\eta}\,\mathbb{E}\left[\sum_{i=1}^{N_{T}}S(H^{z}_{\tau_{i}},v)\right]$
$\quad+\bar{\eta}^{2}\,\mathbb{E}\left[\sum_{i=1}^{N_{T}}S(H^{z}_{\tau_{i}},v)\bigl(1-S(H^{z}_{\tau_{i}},v)\bigr)\right]$
$\quad+\bar{\eta}^{2}\,\mathrm{Var}\left(\sum_{i=1}^{N_{T}}S(H^{z}_{\tau_{i}},v)\right).$
Arguing as in the proof of Proposition 2 yields the formula for $\mathrm{Var}(L^{z}_{T})$ in the statement of the lemma. ∎
In the case $z\equiv 0$ (hence $S(0,v)=v$), the formula of Lemma 1 simplifies to
$\mathrm{Var}(L_{T}^{0})=\mathbb{E}[N_{T}]\big(\sigma^{2}_{\eta}v+\bar{\eta}^{2}v(1-v)\big)+\bar{\eta}^{2}v^{2}\mathrm{Var}(N_{T}).$
This expression can be made fully explicit by using the closed-form expressions for $\mathbb{E}[N_{T}]$ and $\mathrm{Var}(N_{T})$ available for Hawkes processes (see [31, Section 2.1]). For the optimal policy $z^{\ast}$, the quantities appearing in Lemma 1 are computed by Monte Carlo over 100,000100{,}000 samples.
$\mathbb{E}[\eta]$ $\text{Var}(\eta)$ $\sigma(L_{T}^{0})$ $\sigma(L_{T}^{z^{*}})$ $1-[\sigma(L_{T}^{z^{*}})/\sigma(L_{T}^{0})]$
10 10 118.56 51.21 56.80%
10 50 125.04 56.58 54.75%
10 100 132.70 62.65 52.79%
TABLE VI: Standard deviations.
$\mathbb{E}[\eta]$ $\text{Var}(\eta)$ $\pi(L_{T}^{0})$ $\pi(L_{T}^{z^{*}})$ $1-[\pi(L_{T}^{z^{*}})/\pi(L_{T}^{0})]$
10 10 430.55 157.13 63.50%
10 50 432.5 158.74 63.32%
10 100 434.79 160.57 63.07%
TABLE VII: Premia (std deviation principle).
Table VI reports the standard deviations of $L_{T}^{0}$ and $L_{T}^{z^{\ast}}$ for different values of $\sigma^{2}_{\eta}$ (with $\bar{\eta}$ fixed), showing a reduction of roughly $53\%$$57\%$ under optimal prevention. Table VII reports the corresponding premia under (28) with $\theta=0.3$. Across the scenarios considered, the premium under optimal prevention is about $63\%$ lower than in the no-investment benchmark. These results quantify how dynamic optimal cybersecurity investment translates into lower premia.
The values $\pi(L_{T}^{z^{\ast}})$ and $\pi(L_{T}^{0})$ provide natural benchmark lower and upper premia corresponding to optimal prevention and to the absence of additional cybersecurity investment, respectively. More generally, when policyholders are classified into prevention classes using observable proxies (e.g., cybersecurity scores or audits), the model supports premium differentiation across classes: better prevention (corresponding to higher cybersecurity levels) implies lower risk, and therefore lower actuarial premia under standard premium principles. In this sense, the gap between $\pi(L_{T}^{0})$ and $\pi(L_{T}^{z^{\ast}})$ provides a concrete estimate of the premium range induced by prevention under the standard deviation principle.